OpenStudy (anonymous):

What is the solution of the system of equations? 3x+2y=8 x=4 I'm just a little confused on where to start on this. I think after I get the start of it I can do it easily.

OpenStudy (anonymous):

* Solve by graphing

OpenStudy (anonymous):

If x=4 then you can replace x in the first equation with 4 and solve the result for y.

OpenStudy (anonymous):

Is (4,2) correct?

OpenStudy (anonymous):

3x+2y=8 12+2y=8 2y=8-12 y=-4/2